OpenStudy (anonymous):

A doctor recommends that his patients consume at least 1,200 milligrams of calcium each day. A serving of cottage cheese has 75 milligrams of calcium and a serving of milk has 300 milligrams of calcium. If c represents the number of servings of cottage cheese, and m represents the number of servings of milk, which equation below can be used to determine the number of servings of cottage cheese and milk that would guarantee 1,200 milligrams of calcium?

OpenStudy (anonymous):

75m + 300c = 1,200 (c + m)(75 + 300) = 1,200 c + m = 1,200 75c + 300m = 1,200